OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for 15 x sqrt(3) / sqrt(75x^2)

OpenStudy (lgbasallote):

\[\frac{15x \sqrt 3}{\sqrt{75x^2}} \implies \frac{15x \sqrt 3}{\sqrt{25 \times x^2 \times 3}} \implies \frac{15x\sqrt 3}{5x\sqrt 3}\] does that help?

OpenStudy (anonymous):

yeahh

OpenStudy (lgbasallote):

wonderful